# Virtual Sensing for Real-Time Degradation Monitoring of Nuclear Systems: Leveraging DeepONet for Enhanced Sensing Coverage for Digital Twin-Enabling Technology

### v1

@@ Abstract
Effective real-time monitoring technique is crucial for detecting material degradation and maintaining the structural integrity of nuclear systems to ensure both safety and operational efficiency. Traditional physical sensor systems face limitations such as installation challenges, high costs, and difficulties in measuring critical parameters in hard-to-reach or harsh environments, often resulting in incomplete data coverage. Machine learning-driven virtual sensors offer a promising solution by enhancing physical sensor capabilities to monitor critical degradation indicators like pressure, velocity, and turbulence. However, conventional machine learning models struggle with real-time monitoring due to the high-dimensional nature of reactor data and the need for frequent retraining. This paper explores the use of Deep Operator Networks (DeepONet) within a digital twin (DT) framework to predict key thermal-hydraulic parameters in the hot leg of an AP-1000 Pressurized Water Reactor (PWR). In this study, DeepONet is trained with different operational conditions, which relaxes the requirement of continuous retraining, making it suitable for online and real-time prediction components for DT. Our results show that DeepONet achieves accurate predictions with low mean squared error and relative L2 error and can make predictions on unknown data 160,000 times faster than traditional finite element (FE) simulations. This speed and accuracy make DeepONet a powerful tool for tracking conditions that contribute to material degradation in real-time, enhancing reactor safety and longevity.
@@ 1 Introduction
Proper monitoring and inspection of in-service components in nuclear reactors is essential for long-term safety and efficiency, as these components are continuously subjected to extreme temperatures, pressures, and radiation. Among these components, the primary circuit is particularly important, as it removes the immense heat generated in the reactor core, acting as the central component of reactor cooling system. This loop carries highly pressurized water at high velocities, creating significant mechanical stresses on the piping system Zou and Derreberry (2020); Keller and Bilanin (2002). Turbulent coolant flow, especially in areas with bends and joints, leads to localized disturbances Tunstall and Harvey (1968); Dutta et al. (2016); Shimizu et al. (1992) that can induce material degradation mechanisms such as erosion, fatigue, and stress corrosion cracking. Given the critical role of the primary coolant loop, real-time monitoring is essential for detecting early signs of degradation and preventing failures. Tracking key parameters—pressure, velocity, and turbulence—allows operators to identify anomalies such as flow reductions, vibrations, and pressure drops. These insights provide early warnings, often before structural damage occurs, ensuring safe and efficient reactor operation while reducing risks associated with material degradation.
Traditionally, nuclear reactors employ non-destructive testing (NDT) methods, including eddy-current, ultrasonic, radiographic, and visual inspections, to detect degradation in the piping system Khan et al. (2023). These methods are effective at identifying various forms of material degradation, including cracks, corrosion, and fatigue, without causing damage to the system itself. However, these processes are usually carried out through scheduled outages, leading to significant revenue losses. In addition, nuclear piping systems often extend over several miles, making it challenging to inspect the entire system within the limited outage period Sandhu et al. (2024). As a result, while traditional NDT methods remain valuable, they present challenges in terms of achieving comprehensive system coverage and addressing material degradation during routine operations. Real-time monitoring of coolant flow offers an alternative to these scheduled inspections. If during real time monitoring, anomalies or significant deviations from normal parameters are observed, inspections can be scheduled promptly to address specific concerns. This approach not only enhances the safety and reliability of the reactor but also minimizes the need for extensive and costly scheduled outages, as regular scheduled inspections can be replaced by conditional inspections only when there is a clear indication of potential problems Han et al. (2022).
Current methods for real-time coolant flow monitoring in nuclear reactors include ultrasonic, electromagnetic, thermal dispersion, vortex, and differential pressure flow meters Davydov et al. (2021); Tokarz (1983). Additionally, sensors for temperature, pressure, radiation, and chemical composition are placed throughout the coolant piping system to ensure continuous monitoring Kuroze et al. (1996); Lister and Uchida (2015); Jensen et al. (1996). These sensors are typically located at the inlet and outlet of the reactor core, steam generators, and other components U.S. Nuclear Regulatory Commission (2011). When placing sensors, factors such as radiation, temperature, and pressure must be carefully considered to ensure optimal performance and minimize sensor degradation Hossain et al. (2024); Karnik et al. (2024). While the current sensor network for monitoring coolant flow in nuclear reactors provides extensive coverage, ensuring the reactors function properly, there are still a lot of opportunities for further improvement. The main issue with relying solely on physical sensors is the high installation, calibration, and maintenance costs, along with unmonitored areas that create potential blind spots for undetected anomalies. Ideally, sensors would be placed at regular intervals along the piping system for comprehensive monitoring, but this is impractical due to flow disturbances and logistical challenges. While placing sensors at the inlet and outlet of piping segments is feasible, installing them within pipe sections is difficult. Even when sensors are installed and calibrated, failures can complicate detection and replacement, adding operational challenges of relying on extensive physical sensor networks.
To address these limitations, Distributed Fiber Optic (DFO) sensing has emerged as a promising alternative, having been explored for decades in nuclear plants to monitor temperature, flow, and pressure Wu et al. (2021); Ferdinand et al. (2000). Another innovative solution is the use of virtual sensors—software-based models that estimate physical quantities using data and simulations rather than direct physical measurements Zhao et al. (2024a, b); Niresi et al. (2024). These virtual sensors mimic the behavior of physical sensors, providing readings without requiring the installation of actual hardware Sun et al. (2017); Masti et al. (2021); Ofner et al. (2023). They are especially useful in situations where physical placement is not feasible because of limited space or unfavorable weather Ahmed et al. (2013). When used in conjunction with other sensors that are keeping an eye on the same conditions, virtual sensors improve data reliability by reducing signal interference Liu et al. . Virtual sensors are easily adjustable to meet changing needs, unlike real sensors that require actual modifications Barmaz (2023). The concept of virtual sensors in the nuclear sector is relatively new but not entirely unfamiliar. Sevilla et al. Sevilla and Pulido (1998) demonstrate how neural networks can estimate variables in pressurized water reactors by optimizing input selection and network architecture. Ahmed et al. Ahmed et al. (2013) develop virtual sensor networks for accident monitoring in nuclear power plants using statistical relationships among process parameters. Tipireddy et al. Tipireddy et al. (2017) present virtual sensors with Gaussian process models to temporarily replace faulty sensors in nuclear power plants, ensuring accurate measurements and reducing unscheduled shutdowns.
When it comes to monitoring and maintenance of complex in-service systems, Digital Twin technology has gained significant popularity over the past few decades. Digital Twins are virtual models of physical assets that enable real-time monitoring, control, and prediction of system behavior. A typical Digital Twin framework consists of five key components Kobayashi and Alam (2024a, b); Kobayashi et al. (2024a); Kabir and Alam (2010): (1) a prediction module for real-time monitoring, (2) a real-time update module for synchronization, (3) data and signal processing, (4) visualization, and (5) decision-making modules. Within this framework, virtual sensors play a critical role in ensuring the performance and flexibility of the Digital Twin. These sensors continuously collect, analyze, and feed data into the digital twin in real time, supporting both the prediction and data processing components. When establishing digital twins Kobayashi et al. (2024b), virtual sensors can monitor and control physical sensors, offering cost-effective and scalable solutions. This is particularly useful in multi-criteria decision-making applications Martin et al. (2021); Kumar et al. (2022). This capability allows the Digital Twin to maintain an accurate and reliable system representation, even when physical sensors degrade or experience faults. Ultimately, virtual sensors are essential to the scalability of Digital Twin technology, offering a cost-effective and adaptable solution for continuous real-time monitoring, informed decision-making, and system optimization in complex environments.
@@
Emerging neural network technologies, such as the Deep Operator Network (DeepONet) Lu et al. (2021), Fourier Neural Operators (FNO) Li et al. (2020), and Physics-Informed Neural Networks (PINN) Raissi et al. (2019), provide powerful tools for developing virtual sensors that can predict the behavior of complex systems. These models are particularly adept at solving partial differential equations (PDEs) by approximating nonlinear operators and learning mappings between functional spaces. This makes them highly suitable for accurately modeling and predicting the behavior of complex systems with intricate interdependencies, such as those found in nuclear reactors. Among these, the Deep Operator Network (DeepONet) has demonstrated greater efficiency compared to the others. L. Lu, X. Meng and S. Cai et al. in their paper Lu et al. (2022) demonstrate that DeepONet is more flexible than FNO in terms of problem settings and datasets. With PINN as well, we face inconvenience as each new input parameter value requires a separate simulation or retraining. Moreover, PINN struggles to approximate PDEs with strong non-linearity commonly found in practical fluid flow problems. In contrast, DeepONet alleviates the need for retraining by learning operators that map entire functions, rather than specific input-output pairs, across different conditions. Once trained, DeepONet can generalize to a wide range of new inputs without requiring retraining for each new scenario, as it understands the underlying functional relationships between inputs and outputs. This is particularly advantageous for real-time applications, as it can handle varying conditions without the computational burden of frequent retraining. DeepONet approximates linear and nonlinear PDE solution operators using parametric functions as inputs and maps them to various PDE solution functions in output spaces, thereby eliminating the need for these repetitive training processes Koric and Abueidda (2023).
@@
A well-trained neural operator offers the computational efficiency needed for real-time or near real-time predictions, which is crucial for control system optimization Kobayashi et al. (2024a); He et al. (2024). These advantages make DeepONet our preferred choice for its flexibility, scalability, and efficiency. Unlike traditional finite element (FE) simulations, which can be computationally intensive and time-consuming, DeepONet can generate predictions orders of magnitude faster, making it ideal for real-time applications. By training on data from high-fidelity simulations and experimental measurements, DeepONet can accurately predict coolant flow behavior under various operating conditions, which is central to our goal of developing an effective virtual sensor framework for nuclear reactors. The key advantage of using DeepONet for our problem is its trunk network, which processes spatial coordinates to evaluate the output function Lu et al. (2021); He et al. (2024). This network enables us to predict parameters at various points along the pipe based on input data from a single location, allowing for comprehensive condition inference throughout the entire pipe from just one input plane. Combined with near real-time computational ability, DeepONet acts effectively as virtual sensors, providing real-time estimates of flow velocity, turbulence, pressure, and temperature at various points in the primary circuit without physical sensors. Thus, the application of DeepONet offers a solution to many of the challenges posed by current methods for monitoring coolant flow in nuclear reactors.
@@
Figure 1: Schematic diagram of AP-1000 reactor coolant system Namgung and Giang (2016a).
@@
In this work we have studied hot leg as the representation of primary coolant system to demonstrate the effectiveness of DeepONet in predicting the parameters. The hot leg, comprising horizontal and vertical conduits along with elbows, transports highly pressurized, high-temperature water from the reactor pressure vessel (RPV) to the steam generator (SG). Any leakage in the hot leg can lead to serious consequences, including a loss of coolant accident (LOCA), which may impair the reactor’s cooling capacity and pose radiological risks. Thus, it is beneficial to monitor the coolant condition inside the hot leg to ensure proper heat transfer and safe, efficient operation of reactor. However, real time monitoring of key parameters like pressure, velocity, and turbulence kinetic energy inside the hot leg poses challenges due to its critical operating conditions. We aim to develop a real-time monitoring model of thermal-hydraulic conditions within the hot leg of PWR. We have specifically considered the coolant system of AP1000 shown in Figure 1. For this study, we propose to utilize the coolant inlet velocity, which is already monitored through existing sensor networks. We assume that the coolant flow through the hot leg follows the normal operational conditions of a PWR coolant system. Based on this information, we aim to develop a model that predicts specific thermal-hydraulic parameters (pressure, velocity, and turbulence kinetic energy) in the central plane of the hot leg. Monitoring these parameters is important not only for understanding the overall heat transfer performance but also for ensuring the structural integrity of the piping system. High or fluctuating pressure can exert mechanical stress on the piping material, leading to issues such as wear, fatigue, and cracking. Sudden pressure drops may also signal blockages or leaks, which could further accelerate degradation. Monitoring velocity is equally important, as high coolant flow speeds, particularly in areas where the flow direction changes, can induce flow-induced vibrations, leading to mechanical fatigue and increasing the risk of erosion and corrosion. In turn, these conditions can thin the pipe walls over time. Turbulence, especially near bends and joints, amplifies stress on the piping structure, contributing to fatigue and accelerating the erosion-corrosion process. It also introduces hydraulic shocks and uneven temperature distributions, which weaken material integrity over time. Given the critical role that pressure, velocity, and turbulence play in the degradation of piping material, our goal is to develop a real-time monitoring model of the thermal-hydraulic conditions within the hot leg of a PWR.
@@
The key contribution of our work is demonstrating the feasibility and benefits of using DeepONet for real-time monitoring of coolant flow in the primary circuit of a nuclear reactor. Our study highlights the potential of virtual sensors to significantly improve the monitoring and inspection of reactor components. Through comprehensive simulations and experimental validation, we show how this approach delivers highly accurate and timely data, facilitating proactive maintenance and optimizing reactor operations.

### v2

# Virtual Sensing-Enabled Digital Twin Framework for Real-Time Monitoring of Nuclear Systems Leveraging Deep Neural Operators

@@ Abstract
Effective real-time monitoring is a foundation of digital twin technology, crucial for detecting material degradation and maintaining the structural integrity of nuclear systems to ensure both safety and operational efficiency. Traditional physical sensor systems face limitations such as installation challenges, high costs, and difficulty measuring critical parameters in hard-to-reach or harsh environments, often resulting in incomplete data coverage. Machine learning-driven virtual sensors, integrated within a digital twin framework, offer a transformative solution by enhancing physical sensor capabilities to monitor critical degradation indicators like pressure, velocity, and turbulence. However, conventional machine learning models struggle with real-time monitoring due to the high-dimensional nature of reactor data and the need for frequent retraining. This paper introduces the use of Deep Operator Networks (DeepONet) as a core component of a digital twin framework to predict key thermal-hydraulic parameters in the hot leg of an AP-1000 Pressurized Water Reactor (PWR). DeepONet serves as a dynamic and scalable virtual sensor by accurately mapping the interplay between operational input parameters and spatially distributed system behaviors. In this study, DeepONet is trained with different operational conditions, which relaxes the requirement of continuous retraining, making it suitable for online and real-time prediction components for digital twin. Our results show that DeepONet achieves accurate predictions with low mean squared error and relative L2 error and can make predictions on unknown data 1400 times faster than traditional CFD simulations. This speed and accuracy enable DeepONet to synchronize with the physical system in real-time, functioning as a dynamic virtual sensor that tracks degradation-contributing conditions. By providing comprehensive system insights and predictive capabilities, DeepONet firmly establishes itself as a critical tool within the Digital Twin framework for nuclear systems.
@@ 1 Introduction
Proper monitoring and inspection of in-service components in nuclear reactors is essential for long-term safety and efficiency, as these components are continuously subjected to extreme temperatures, pressures, and radiation. Among these components, the primary circuit is particularly important, as it removes the immense heat generated in the reactor core, acting as the central component of the cooling system. This loop carries highly pressurized water at high velocities, creating significant mechanical stresses on the piping system Zou and Derreberry (2020); Keller and Bilanin (2002). Turbulent coolant flow, especially in areas with bends and joints, leads to localized disturbances Tunstall and Harvey (1968); Dutta et al. (2016); Shimizu et al. (1992) that can induce material degradation mechanisms such as erosion, fatigue, and stress corrosion cracking. Given the significant role of the primary coolant loop, real-time monitoring is essential for detecting early signs of degradation and preventing failures. Tracking key parameters such as pressure, velocity, and turbulence, enables the detection of deviations from normal operational conditions, such as flow reductions, vibrations, or pressure drops, which can serve as early indicators of potential material degradation. These insights provide early warnings, often before structural damage occurs, ensuring safe and efficient reactor operation while reducing risks associated with material degradation.
Non-destructive testing (NDT) methods, such as eddy-current, ultrasonic, radiographic, and visual inspections, are considered standard practices for degradation monitoring in nuclear reactors Khan et al. (2023). These methods effectively identify various forms of material degradation, including cracks, corrosion, and fatigue, without damaging the system itself. However, they are usually performed during scheduled outages, resulting in significant revenue losses. Additionally, nuclear piping systems often extend over several miles, making it challenging to inspect the entire system within a limited outage period Sandhu et al. (2024). In-line inspections (ILI) are widely recognized as a standard approach for assessing internal pipeline conditions, including wall thinning, corrosion, and structural integrity Kishawy and Gabbar (2010); Larsson (2017). However, they are also conducted periodically and has similar issue as NDT methods. Thus, while traditional periodic inspection methods remain valuable, they present challenges in addressing material degradation during routine operations. Continuous monitoring of coolant flow in real time offers a promising alternative to scheduled inspections. By identifying anomalies in operational parameters as they occur, real time monitoring enables prompt scheduling of inspections to address specific concerns before they escalate into critical issues. Current pipeline monitoring systems use advanced simulation models to enable real-time monitoring of pressure conditions. These models offer valuable insights into operational parameters, particularly in areas where direct measurements are either impractical or impossible Larsson (2017). However, simulation models take a long time to get results which make them unsuitable for continuous real-time monitoring tool. Real-time monitoring can address the shortcomings of periodic inspections if it can provide uninterrupted oversight of system parameters.
This uninterrupted view of entire reactor can be achieved through Digital Twin technology. A Digital Twin creates a virtual replica of the physical system, allowing for real-time monitoring, control, and prediction of system behavior Kobayashi and Alam (2024a, b); Kobayashi et al. (2024a). By dynamically updating itself with real-time data, a Digital Twin provides a comprehensive and synchronized view of the reactor’s operational state. This capability enables proactive maintenance, fault detection, and optimization of reactor operations, reducing the reliance on periodic inspection and minimizing operational risks. By continuously integrating operational data, Digital Twin frameworks enable predictive modeling of degradation processes, such as material fatigue and erosion-corrosion, ensuring timely interventions. Digital Twin framework rely heavily on sensors as they work as a connection between the physical reactor and its digital counterparts Hossain et al. (2024a). Sensors for coolant flow monitoring in nuclear reactors include ultrasonic, electromagnetic, and thermal dispersion flow meters Davydov et al. (2021); Tokarz (1983), as well as sensors for key parameters like temperature, pressure, and radiation Kuroze et al. (1996); Lister and Uchida (2015); Jensen et al. (1996). These sensors are strategically placed at critical locations, such as the inlet and outlet of the reactor core, steam generators, and other components, to ensure continuous monitoring of system conditions U.S. Nuclear Regulatory Commission (2011), as well as to preserve sensor health and minimize degradation Hossain et al. (2024b); Karnik et al. (2024). However, for a successful Digital Twin framework, comprehensive coverage of the entire reactor system is needed, which cannot be achieved solely through physical sensors. While placing sensors at the inlet and outlet of piping segments is feasible, installing them within pipe sections is impractical as it will cause flow disturbances and also has logistical challenges. Though DFO can address coverage limitations, it is very expensive and cannot be implemented in existing reactors.
Virtual sensors can address these limitations of physical sensors. They are software-based models that estimate physical quantities using data and simulations rather than direct physical measurements Zhao et al. (2024a, b); Niresi et al. (2024). They mimic the behavior of physical sensors, providing readings without requiring the installation of actual hardware Sun et al. (2017); Masti et al. (2021); Ofner et al. (2023). Integrated within a Digital Twin, virtual sensors utilize data from sporadically placed physical sensors to predict values in unseen or unmonitored areas, providing critical insights into system conditions in real time and offering a complete view of the reactor’s operational state. Unlike physical sensors, virtual sensor-based Digital Twin frameworks are not constrained by installation or environmental limitations, making them ideal for monitoring hard-to-reach or harsh reactor environments Ahmed et al. (2013). Furthermore, Digital Twin frameworks utilize existing sensor networks to operate virtual sensors, reducing the need for modifications to established reactor designs while still expanding monitoring capabilities. They also improve data reliability by reducing signal interference and are easily adjustable to accommodate changes in system requirements Liu et al. ; Barmaz (2023). This adaptability ensures the Digital Twin remains accurate and reliable, even when physical sensors degrade or fail.
While virtual sensing is still a relatively new concept in nuclear applications, its potential to enhance system reliability and efficiency has been demonstrated over past years. For instance, Sevilla et al. Sevilla and Pulido (1998) showed how neural networks could estimate variables in pressurized water reactors by optimizing input selection and network architecture. Ahmed et al. Ahmed et al. (2013) developed virtual sensor networks for accident monitoring in nuclear plants, while Tipireddy et al. Tipireddy et al. (2017) introduced Gaussian process-based virtual sensors to replace faulty physical sensors, reducing unscheduled downtime. These examples underscore the critical role virtual sensing plays in Digital Twin frameworks, enabling real-time monitoring, fault tolerance, and system optimization in nuclear reactors.
@@
Emerging neural network technologies, such as the Deep Operator Network (DeepONet) Lu et al. (2021), Fourier Neural Operators (FNO) Li et al. (2020), and Physics-Informed Neural Networks (PINN) Raissi et al. (2019) provide powerful tools for developing virtual sensors. These models are adept at solving partial differential equations (PDEs) by approximating nonlinear operators and learning mappings between functional spaces. This makes them highly suitable for modeling and predicting the behavior of complex, interdependent systems, such as nuclear reactors. Among these algorithms, the Deep Operator Network (DeepONet) has demonstrated greater efficiency compared to others. Lu et al. in their recent paper Lu et al. (2022) show that, DeepONet is more flexible than FNO in terms of problem settings and datasets. With PINN, as well, each new parameter requires separate simulations or retraining. Moreover, PINN struggles to approximate PDEs with strong non-linearity commonly found in practical fluid flow problems. In contrast, DeepONet alleviates the need for retraining by learning operators that map entire functions, rather than specific input-output pairs, across different conditions. Once trained, DeepONet can generalize to a wide range of new inputs without requiring retraining for each new scenario, as it understands the underlying functional relationships between inputs and outputs. This is advantageous for real-time applications, adapting to varying conditions without frequent retraining. DeepONet approximates linear and nonlinear PDE solution operators using parametric functions as inputs, mapping them to output spaces, thereby eliminating retraining requirements Koric and Abueidda (2023).
@@
A well-trained neural operator offers the computational efficiency needed for real-time or near real-time predictions, which is crucial for control system optimization Kobayashi et al. (2024a); He et al. (2024). These advantages make DeepONet our preferred algorithm for establishing Digital Twin. Unlike traditional finite element/volume (FEM/FVM) simulations, which are computationally intensive and time-consuming, DeepONet generates predictions orders of magnitude faster, making it ideal for real-time applications. By training on high-fidelity simulations and experimental data, DeepONet accurately predicts coolant flow behavior under various operating conditions, which is central to developing a virtual sensor framework for nuclear reactors. The key advantage of using DeepONet for our problem is its trunk network, which processes spatial coordinates to evaluate the output function Lu et al. (2021); He et al. (2024). This network predicts parameters at various pipe locations based on input data from a single location, providing comprehensive pipe condition inference. Combined with near real-time capability, DeepONet provides highly accurate estimates of flow velocity, turbulence, pressure, and temperature at key primary circuit locations without physical sensors.
@@
Figure 1: Schematic diagram of AP-1000 reactor coolant system Namgung and Giang (2016).
@@
In this work, we study the hot leg as the representation of the primary coolant system to demonstrate DeepONet’s effectiveness. The hot leg, consisting of horizontal and vertical conduits with elbows, transports high-pressure, high-temperature water from the reactor pressure vessel (RPV) to the steam generator (SG). Leakage in the hot leg can cause serious issues, including a loss of coolant accident (LOCA), impairing reactor’s cooling capacity and posing radiological risks. Monitoring coolant conditions in the hot leg is crucial to ensure proper heat transfer and reactor safety. However, real-time monitoring of key parameters like pressure, velocity, and turbulence kinetic energy inside the hot leg is challenging due to critical operating conditions. We focus on developing a real-time monitoring model of thermal-hydraulic conditions for the hot leg of the AP1000, as shown in Figure 1. We use the coolant inlet velocity, already monitored through existing sensors, assuming normal PWR operational conditions. Based on this information, we develop a model to predict thermal-hydraulic parameters (pressure, velocity, and turbulence kinetic energy) in the hot leg’s central plane. Monitoring these parameters is important for assessing heat transfer performance and ensuring structural integrity. For instance, high or fluctuating pressure stresses piping, causing wear, fatigue, and cracking, while sudden drops in pressure may signal blockages or leaks, accelerating degradation. Monitoring velocity is equally important, as high coolant flow velocity, particularly in areas where the flow direction changes, can cause flow-induced vibrations. These vibrations lead to mechanical fatigue and increase the risk of erosion and corrosion, gradually thinning the pipe walls. Similarly, turbulence, especially near bends and joints, amplifies stress on the piping structure, accelerating fatigue and the erosion-corrosion process. It also introduces hydraulic shocks and uneven temperature distributions, further weakening material integrity over time. Monitoring these parameters in real time enables operators to correlate abnormal flow patterns with potential degradation, allowing for timely maintenance decisions. Given their role in degradation, our goal is to develop a real-time thermal-hydraulic monitoring model for the hot leg of a PWR.
@@
The key contribution of our work is demonstrating the feasibility and benefits of using DeepONet within a Digital Twin framework for real-time monitoring of coolant flow in the primary circuit of a nuclear reactor. By integrating virtual sensors, our study highlights the potential of Digital Twins to significantly improve the monitoring and inspection of reactor components. Through comprehensive simulations and experimental validation, we show how this approach delivers accurate and timely data, enabling the detection of operational conditions indicative of degradation and facilitating proactive maintenance and optimizing reactor operations.